OpenStudy (anonymous):

26.4 is what percent of 66

OpenStudy (ivettef365):

what is 26.4 divided by 66 ?

OpenStudy (anonymous):

is/of = %/100 So 26.4/66 = x/100 Cross Multiply. 26.4(100)=66x 2640=66x 2640/66=x x=37.27%

OpenStudy (ivettef365):

the only thing you need to do is divide 26.4/66 and you get answer nothing else